### v1

# Conformal Risk Minimization with Variance Reduction

Abstract
Conformal prediction (CP) is a distribution-free framework for achieving probabilistic guarantees on black-box models. CP is generally applied to a model post-training. Recent research efforts, on the other hand, have focused on optimizing CP efficiency during training. We formalize this concept as the problem of conformal risk minimization (CRM). In this direction, conformal training (ConfTr) by Stutz et al. (2022) is a technique that seeks to minimize the expected prediction set size of a model by simulating CP in-between training updates. Despite its potential, we identify a strong source of sample inefficiency in ConfTr that leads to overly noisy estimated gradients, introducing training instability and limiting practical use. To address this challenge, we propose variance-reduced conformal training (VR-ConfTr), a CRM method that incorporates a variance reduction technique in the gradient estimation of the ConfTr objective function. Through extensive experiments [^1] We release our code at https://github.com/nooranisima/conformal-risk-minimization-w-variance-reduction-code on various benchmark datasets, we demonstrate that VR-ConfTr consistently achieves faster convergence and smaller prediction sets compared to baselines.
1 Introduction
Consider a classification task with input (features) $X\in\mathcal{X}$ and corresponding label $Y\in\mathcal{Y}=\{1,\ldots,K\}$. In supervised learning, we are interested in approximating the posterior probability $\pi(y|x)=\mathbb{P}\left({Y=y\,|\,X=x}\right)$ by tuning some $\theta$-parameterized family of predictors $\pi_{\theta}(y|x)$ for example, neural network models with a softmax activation at the output layer. Typically, the final label prediction would be $\delta_{\theta}(x)=\operatorname*{arg\,max}_{y\in\mathcal{Y}}\pi_{\theta}(y|x)$, and a common metric for performance is the accuracy, which measures the proportion of testing samples whose predicted label matches the true label. While the accuracy is a key performance metric, in safety-critical applications with a downstream decision maker it is crucial not only to predict accurately but also to quantify the uncertainty associated with a prediction.
Conformal prediction (CP) is a distribution-free, principled framework that is used to provide formal probabilistic guarantees for black-box models (Vovk et al. (2005); Shafer & Vovk (2008); Angelopoulos et al. (2023)), with exemplar applications in computer vision (Angelopoulos et al. (2020)), large language models (Mohri & Hashimoto (2024),Kumar et al. (2023)) and path planning (Lindemann et al. (2023)). Given a model $\pi_{\theta}(y|x)$, CP enables the construction of set predictors $C_{\theta}:\mathcal{X}\to 2^{\mathcal{Y}}$ (where $2^{\mathcal{Y}}$ is the power set of $\mathcal{Y}$) such that the true label is contained in the set of predicted labels with high probability. This can be formalized via the notion of marginal coverage.
Definition 1.1 (Marginal coverage).
We say that a set predictor $C_{\theta}:\mathcal{X}\to 2^{\mathcal{Y}}$ satisfies marginal coverage with miscoverage rate $\alpha\in(0,1)$ if $\mathbb{P}\left({Y\in C_{\theta}(X)}\right)\geq 1-\alpha$.
Marginal coverage can be readily obtained in CP via a process called calibration, which only requires access to a so-called calibration set of data that is statistically exchangeable with the test data. However, one of the main challenges in CP is the efficiency of the prediction sets - namely the size of the sets $C_{\theta}(x)$ often referred to as length efficiency (Fontana et al., 2023). For instance, while it is possible to trivially achieve the desired coverage by including the entire label space in $C_{\theta}(x)$, such an approach results in non-informative and excessively large prediction sets. An efficient $C_{\theta}(x)$ is as small as possible while still maintaining the coverage guarantee.
Various existing approaches, including the works by Romano et al. (2020); Yang & Kuchibhotla (2024); Bai et al. (2022), address the efficiency challenge by refining the conformal prediction procedure applied post-training to a black-box model. These methods, though effective, are constrained by the performance of the pre-trained model $\pi_{\theta}(y|x)$ on which they are applied. On the other hand, recent research efforts (Dheur & Taieb, 2024; Cherian et al., 2024; Einbinder et al., 2022; Stutz et al., 2022; Bellotti, 2021) have focused on integrating CP directly into the training process. This provides a way to improve the CP efficiency also in the model optimization phase - when learning the parameter of the models - enabling a higher degree of control over the probabilistic guarantees efficiency. In this work, we formulate this approach as conformal risk minimization (CRM) and we focus on CRM for length efficiency optimization. We consider a setting similar to  Stutz et al. (2022), who proposed conformal training (ConfTr), an algorithm achieving promising performance in improving the length-efficiency of the prediction sets constructed post-training.
Despite encouraging preliminary results, ConfTr introduces significant optimization challenges, particularly due to the use of differentiable approximations of CP sets. Indeed, ConfTr requires differentiating a loss function obtained simulating CP during training. This, in turn, requires accurately estimating the population quantile of the conformity scores and its gradient, which can be difficult with the limited data available in each mini-batch. Hence, training can exhibit large fluctuations in the loss and slow convergence, thereby reducing the practical applicability of the method.
In this work, we address these challenges by introducing variance-reduced conformal training (VR-ConfTr), a novel CRM algorithm leveraging a variance reduction technique for the estimation of quantiles’ gradients. Relative to confTr, our proposed VR-ConfTr solution significantly stabilizes training - leading to faster convergence, and consistently enhances the length efficiency of post-training conformal prediction sets when compared against baselines.
Figure 1: Comparison of VR-ConfTr, ConfTr, and Baseline model on Fashion-MNIST. VR-ConfTr shows faster convergence with lower training loss and achieves more efficient set sizes during training. Post-training evaluation shows VR-ConfTr has higher accuracy than ConfTr.
1.1 Contributions
Our contributions can be summarized as follows:
Conformal risk minimization. We formulate conformal risk minimization (CRM) as a framework for training a parameterized predictor that learns according to CP efficiency metrics.
A “plug-in” algorithm. Focusing on CRM for length efficiency optimization, we provide a novel analysis for the variance of the ConfTr (Stutz et al., 2022) method, which shows the need for improved estimators of the quantile gradients. Then, we introduce the pipeline of variance-reduced conformal training (VR-ConfTr), our proposed algorithm to overcome this challenge, which leverages a “plug-in” step to incorporate improved estimates of quantiles’ gradients in the training.
Novel variance reduction technique. Building on a fundamental result, which characterizes the gradient of the population quantile as a conditional expectation, we propose a novel estimator for quantile gradients whose variance is provably reduced with the training batch size. This novel estimator can be seamlessly integrated into VR-ConfTr. We analyze the bias-variance trade-off of this novel estimator and establish its precise relationship with the conformity measures associated to a predictor $\pi_{\theta}(y|x)$.
Empirical validations. We extensively analyze our method on various benchmark and real-world datasets, including MNIST, FMNIST, KMNIST and OrganAMNIST. Our results demonstrate that VR-ConfTr consistently and significantly improves the efficiency and stability of conformal training for length efficiency optimization. See Fig. 1 for an example of performance evaluation of VR-ConfTr against ConfTr and a baseline trained using standard cross-entropy loss.
Broad applicability. Our approach and novel variance reduction technique can be integrated into any CRM method that requires quantile gradient estimation, extending its utility to a large class of conformal prediction frameworks and learning models.
Figure 2: In this figure, we illustrate the VR-ConfTr pipeline and position it with respect to a typical CP procedure.
1.2 Related Work
A large body of research has focused on optimizing length-efficiency in CP. We now review some recent literature in this area. We first (i) review approaches that apply CP post-training to black-box models, and then (ii) review the recent efforts in coupling CP and model training, in what we call conformal risk minimization (CRM) approaches. For (i), recent algorithmic developments address improving length efficiency through better conformity score design Romano et al. (2020); Yang & Kuchibhotla (2024); Amoukou & Brunel (2023); Deutschmann et al. (2024); Luo & Zhou (2024). From another perspective Kiyani et al. (2024); Bai et al. (2022); Yang & Kuchibhotla (2021); Colombo & Vovk (2020) focus on designing better calibration procedures. Particularly, Kiyani et al. (2024) propose an optimization technique for the calibration threshold $\tau$. On the other hand, Bai et al. (2022); Yang & Kuchibhotla (2021) formulate CP as a constrained optimization problem, minimizing the average prediction interval width with valid empirical coverage. These efforts do not fall under the CRM framework because they focus on learning low-dimensional hyper-parameters for pre-trained models as opposed to fully guiding the training of the parameters $\theta$ of a model $\pi_{\theta}(y|x)$.
Conformal risk minimization. There is a growing body of work (Einbinder et al., 2022; Cherian et al., 2024; Stutz et al., 2022; Bellotti, 2021; Yan et al., 2024) integrating ideas from conformal prediction in order to directly train a model for improved uncertainty quantification. Cherian et al. (2024) train a score function, rather than a point predictor, subject to conditional coverage constraints (Gibbs & Candes (2021)). Einbinder et al. (2022) utilize conformal prediction insights in order to mitigate overconfidence in multi-class classifiers by minimizing a carefully designed loss function. Stutz et al. (2022) proposed conformal training (ConfTr), in which length efficiency optimization is tackled by defining a loss function obtained by simulating conformal prediction during training. We will extensively describe this approach in the next section. Yan et al. (2024) uses a similar training pipeline to Stutz et al. (2022) in order to minimize the inefficiency of their proposed conformal predictor. The early work by Bellotti (2021) considered an approach analogous to ConfTr in that the authors simulate conformal prediction during training. However, the algorithm provided by Bellotti (2021) treats the quantile-threshold as fixed and not as a function of the model parameters, and it has been extensively shown by Stutz et al. (2022) that this approach provides inferior performance with respect to ConfTr.
2 Problem Formulation
Let us consider a parameterized model of logits $f_{\theta}:\mathcal{X}\to\mathbb{R}^{K}$ and let $\pi_{\theta}(x)=\mathrm{softmax}(f_{\theta}(x))$ denote the corresponding predicted probabilities. The objective of a conformal prediction algorithm is to construct a set predictor $C_{\theta}:\mathcal{X}\to 2^{\mathcal{Y}}$ starting from the model $f_{\theta}$ in such a way that $C_{\theta}$ achieves marginal coverage. One common way to achieve this is via a thresholding (THR) set predictor (Vovk et al. (2005)), $C_{\theta}(x;\tau)=\{y\in\mathcal{Y}:E_{\theta}(x,y)\geq\tau\}$ for some well chosen threshold $\tau$ and conformity score $E_{\theta}(x,y)$, which can be any heuristic notion of uncertainty regarding label $y$ upon input $x$ for the predictor $f_{\theta}(\cdot)$. Some choices for the conformity score include (i) the predicted probabilities $E_{\theta}(x,y)=\pi_{\theta}(y|x)=[\pi_{\theta}(x)]_{y}$, (ii) the logits $E_{\theta}(x,y)=[f_{\theta}(x)]_{y}$, and (iii) the predicted log-probabilities $E_{\theta}(x,y)=\log\pi_{\theta}(y|x)$. If we knew the marginal distribution for $(X,Y)$, then marginal coverage could be readily achieved by setting $\tau=\tau(\theta)=Q_{\alpha}(E_{\theta}(X,Y))$ where $Q_{\alpha}$ denotes the population quantile of some scalar random variable. Indeed,
$\mathbb{P}\left({Y\in C_{\theta}(X;\tau}\right))$ $=\mathbb{P}\left({E_{\theta}(X,Y)\geq\tau}\right)\geq 1-\alpha$ (1)
directly from $\tau=Q_{\alpha}(E_{\theta}(X,Y))$. In practice, however, the marginal distribution of $(X,Y)$ is not known. Instead, we can estimate $\tau(\theta)=Q_{\alpha}(E_{\theta}(X,Y))$ from data that, as long as it satisfies the principle of exchangeability, can be used to ascertain marginal coverage of $C_{\theta}(x;\tau)$.
2.1 Conformal Risk Minimization
As we outlined in the introduction, recent research efforts have attempted to combine training and conformal prediction (CP) into one, as opposed to using CP only as a post-training method. Here, we formalize this by borrowing terminology from statistical supervised learning and by introducing the problem of conformal risk minimization (CRM). CRM can be understood as a framework for training a parameterized predictor that learns according to some CP efficiency metric, and can be formulated as follows:
$\min_{\theta\in\Theta}L(\theta):=\mathbb{E}\left[{\ell(C_{\theta}(X),Y)}\right]$ (CRM)
for some conformal loss $\ell$, where $C_{\theta}(x)$ is a conformalized predictor. This problem is closely related to the conformal risk control explored byAngelopoulos et al. (2022).
2.2 ConfTr (Stutz et al., 2022)
Stutz et al. (2022) introduced conformal training (ConfTr), which we can categorize as a CRM approach for length efficiency optimization. In particular, ConfTr focuses on reducing inefficiency of calibrated classifiers, quantified by the target size of predicted sets. This can be understood as the problem in (CRM) with $\ell(C,y)=\max(0,|C|-\kappa)$ for some target size $\kappa$ (intended to discourage no predictions at all). In this regard, it is worth noting that the earlier work of Sadinle et al. (2019) was the first to study the closely related problem of least ambiguous set-valued classifiers, which corresponds to $l(C,y)=|C|$. An important aspect of the work of Stutz et al. (2022) is that the authors relaxed the CRM problem with target size conformal loss $\ell(C,y)=\max(0,|C|-\kappa)$ into a smooth one in $\theta$, in order to allow gradient-based optimization to be employed. In particular, the authors relax the prediction set $C_{\theta}(x;\tau)$ into the smooth prediction “set” (vector) $\mathbf{C}_{\theta}(x;\tau)\in[0,1]^{K}$ with relaxed binary indicator variables, given by
$[\mathbf{C}_{\theta}(x;\tau)]_{y}=\sigma\left(\frac{E_{\theta}(x,y)-\tau}{T}\right)$ (2)
for $y\in\mathcal{Y}$, where $\sigma(\cdot)$ denotes the sigmoid function and $T>0$ a “temperature” hyper-parameter such that $[\mathbf{C}_{\theta}(x;\tau)]_{y}\to 1_{E_{\theta}(x,y)\geq\tau}$ as $T\to 0$, with $1_{A}$ the indicator function for condition $A$. Further, the prediction set size $|C_{\theta}(x;\tau)|$ is relaxed into the smooth prediction set size $\sum_{y\in\mathcal{Y}}[\mathbf{C}_{\theta}(x;\tau)]_{y}=\mathbf{1}_{K}^{\mathsf{T}}\mathbf{C}_{\theta}(x;\tau)$. With this, the problem solved by Stutz et al. (2022) can be written as
$\min_{\theta\in\Theta}L(\theta)=\log\mathbb{E}\left[{\Omega\big{(}\mathbf{C}_{\theta}(X;\tau(\theta)\big{)}}\right]$ (3)
with $\Omega(\mathbf{C})=\max(0,\mathbf{1}_{K}^{\mathsf{T}}\mathbf{C}-\kappa)$. Additionally, the authors explored other terms, such as a configurable class-conditional “coverage loss”
$\mathcal{L}(\mathbf{C},y)=\sum_{y^{\prime}\in\mathcal{Y}}[\mathbf{L}]_{yy^{\prime}}\big{(}(1-[\mathbf{C}]_{y^{\prime}}\delta_{yy^{\prime}}+[\mathbf{C}]_{y^{\prime}}(1-\delta_{yy^{\prime}})\big{)},$
as well as a possible base loss (such as cross entropy) and regularizer. The $\log$ term in (3) is used for numerical stability reasons for the gradient-based optimizers employed by the authors. Let us abstract these factors into the problem
$\min_{\theta\in\Theta}L(\theta):=h(\mathbb{E}\left[{\ell(\theta,\tau(\theta),X,Y)}\right])+R(\theta)$ (ConfTr-risk)
to be solved via a gradient-based method for some monotone transformation $h(\cdot)$, conformal loss $\ell(\cdot)$, and regularizer $R(\cdot)$. The underlying assumption, just as in any supervised learning task, is that the marginal distribution of $(X,Y)$ is unknown but that instead we can collect some i.i.d. training data $\mathcal{D}=\{(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})\}$. With this, an issue presents itself in that, unlike a typical loss function, we cannot evaluate $\frac{\partial}{\partial\theta}[\ell(\theta,\tau(\theta),X_{i},Y_{i})]$ from knowledge alone of $\theta,X_{i},Y_{i}$, because $\tau(\theta)=Q_{\alpha}(E_{\theta}(X,Y))$ is a function of the distribution of $(X,Y)$ and not a mere transformation. To resolve this issue, Stutz et al. (2022) propose their ConfTr algorithm, which randomly splits a given batch $B$ into two parts, which they refer to as calibration batch $B_{\mathrm{cal}}$ and prediction batch $B_{\mathrm{pred}}$. With this, the authors advocate for employing any smooth (differentiable) quantile estimator algorithm for $\tau(\theta)$ using the calibration batch. Then, they propose using this estimator to compute a sampled approximation of (ConfTr-risk), replacing expectations by sample means constructed using the prediction batch. Let $\hat{L}(\theta)$ denote the end-to-end empirical approximation of $L(\theta)$ in terms of the current parameters $\theta$. Once $\hat{L}(\theta)$ is constructed, the authors advocate for a (naive) risk minimization procedure where $\frac{\partial\hat{L}}{\partial\theta}(\theta)$ is computed and passed to an optimizer of choice.
2.3 Variance Analysis for ConfTr
Consider i.i.d. samples $\{(X_{i},Y_{i})\}_{i=1}^{n}$ from which we seek to estimate $\tau(\theta)=Q_{\alpha}(E_{\theta}(X,Y))$ and let $E_{(1)}(\theta)\leq\ldots\leq E_{(n)}(\theta)$ denote the order statistics corresponding to the scalar random variables $E_{\theta}(X_{1},Y_{1}),\ldots,E_{\theta}(X_{n},Y_{n})$. Unlike the expectation and covariance matrix of a random vector, there is no universal consensus on an estimator for the population quantile of scalar random variables. Hyndman & Fan (1996) summarized and unified a significant portion of the various estimators proposed in the literature at the time. Following the notation of the aforementioned work, we will consider estimators of the form
$\hat{\tau}(\theta)=\gamma E_{(j)}(\theta)+(1-\gamma)E_{(j+1)}(\theta)$ (4)
for some $\gamma=\gamma(j,g)\in[0,1]$ with $j=\lfloor\alpha n+r\rfloor$ and $g=\alpha n+r-j$, where $r\in\mathbb{R}$ is a hyper-parameter. Other estimators have been proposed since, and even at the time of (Hyndman & Fan, 1996). However, the majority of statistical packages implement, by default, an estimator of the form (4). Other approaches have been proposed in the literature, via kernel-based methods, variational methods, and dispersion-based methods. With this estimator of equation (4), we see that
$\frac{\partial\hat{\tau}}{\partial\theta}(\theta)=\gamma\frac{\partial E_{(j)}}{\partial\theta}(\theta)+(1-\gamma)\frac{\partial E_{(j+1)}}{\partial\theta}(\theta).$ (5)
Further, if $\omega(\theta):[n]\to[n]$ denotes the permutation of indices $[n]:=\{1,\ldots,n\}$ that correspond to the order statistics, i.e. $E_{(j)}(\theta)=E_{\theta}(X_{\omega_{j}(\theta)},Y_{\omega_{j}(\theta)})$ with $\omega(\theta)=(\omega_{1}(\theta),\ldots,\omega_{n}(\theta))$, we see that $\omega(\theta)$ is piecewise constant (or approximately so when using a smooth sorting such as in (Blondel et al., 2020; Cuturi et al., 2019)), and thus $\frac{\partial\omega}{\partial\theta}(\theta)\approx 0$. By the chain rule, it follows that $\frac{\partial E_{(j)}}{\partial\theta}(\theta)\approx\frac{\partial E}{\partial\theta}(\theta,X_{\omega_{j}(\theta)},Y_{\omega_{j}(\theta)})$, where $E(\theta,X,Y)=E_{\theta}(X,Y)$. Since $E_{(j)}(\theta)\approx\tau(\theta)$ and $E_{(j+1)}(\theta)\approx\tau(\theta)$, and noting that the samples $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ are i.i.d., then $(X_{\omega_{j}(\theta)},Y_{\omega_{j}(\theta)})$ and $(X_{\omega_{j+1}(\theta)},Y_{\omega_{j+1}(\theta)})$ are approximately independent and approximately distributed as equal to the distribution of $(X,Y)$ when conditioned on $E_{\theta}(X,Y)=\tau(\theta)$. Hence,
$\mathbb{E}\left[{\frac{\partial\hat{\tau}}{\partial\theta}(\theta)}\right]$ $\approx\mathbb{E}\left[{\frac{\partial E}{\partial\theta}(\theta,X,Y)\,\Big{|}\,E_{\theta}(X,Y)=\tau(\theta)}\right]$ (6)
$\mathrm{cov}\left({\frac{\partial\hat{\tau}}{\partial\theta}(\theta)}\right)$ $\approx(\gamma^{2}+(1-\gamma)^{2})\,\mathrm{cov}\left({\frac{\partial E}{\partial\theta}(\theta,X,Y)\,\Big{|}\,E_{\theta}(X,Y)=\tau(\theta)}\right).$ (7)
Inspecting (7), we can see that the variance of the naive estimator $\frac{\partial\hat{\tau}}{\partial\theta}(\theta)$ for $\frac{\partial\tau}{\partial\theta}(\theta)$ is approximately constant when the sample size is moderately large. In particular, the variance is approximately $\mathcal{O}(1)$, which is quite sample inefficient as it does not decrease as the sample size increases. On the other hand, by differentiating the conformal training loss, i.e.
$\frac{\partial}{\partial\theta}[\ell(\theta,\hat{\tau}(\theta),x,y)]=\frac{\partial\ell}{\partial\theta}(\theta,\hat{\tau}(\theta),x,y)+\frac{\partial\ell}{\partial\tau}(\theta,\hat{\tau}(\theta),x,y)\frac{\partial\hat{\tau}}{\partial\theta}(\theta),$ (8)
it becomes apparent that poor estimator variance for $\frac{\partial\hat{\tau}}{\partial\theta}(\theta)$ will bottleneck sample efficiency in the estimation of $\frac{\partial L}{\partial\theta}(\theta)$ obtained by replacing $\tau(\theta)$ in (ConfTr-risk) with $\hat{\tau}(\theta)$ and using the prediction batch to approximate the expectations. In the next section, we present our proposed solution to address this issue.
3 Variance-Reduced Conformal Training
In order to surpass the shortcoming of ConfTr described in the previous section, let us first note that the gradient of the conformal risk (ConfTr-risk) can be written as
$\frac{\partial L}{\partial\theta}(\theta)=h^{\prime}(\mathbb{E}\left[{\ell(\theta,\tau(\theta),Z)}\right])\left(\mathbb{E}\left[{\frac{\partial\ell}{\partial\theta}(\theta,\tau(\theta),Z)}\right]+\mathbb{E}\left[{\frac{\partial\ell}{\partial\tau}(\theta,\tau(\theta),Z)}\right]\frac{\partial\tau}{\partial\theta}(\theta)\right),$ (9)
where $h^{\prime}$ denotes the derivative of $h$, $Z=(X,Y)$, and noting that we dropped the regularizer for simplicity. Additionally, we can exploit the following relationship to further characterize ConfTr as well as to design a variance-reduced alternative:
Proposition 3.1 (Quantile Sensitivity (Hong, 2009)).
Suppose that $X$ is absolutely continuous and $E_{\theta}(x,y)$ is continuously differentiable in $\theta$ and $x$. Then, for every $\theta\in\Theta$,
$\frac{\partial\tau}{\partial\theta}(\theta)=\mathbb{E}\Bigg{[}\frac{\partial E}{\partial\theta}(\theta,X,Y)\,\Big{|}\,E_{\theta}(X,Y)=\tau(\theta)\Bigg{]}.$ (10)
In Appendix A, we provide a rigorous proof for the above proposition, which was carried out independently from that of the equivalent result of Hong (2009) (namely, Theorem 2). However, note that the assumptions in (Hong, 2009) are less restrictive than the ones we use. Further, the author explores more deeply the connections between $\tau(\theta)$ and $\frac{\partial\tau}{\partial\theta}(\theta)$.
Equipped with the above proposition, we can compare (6) and (10) to see that, despite the poor sample efficiency of the naive estimator for $\frac{\partial\tau}{\partial\theta}(\theta)$, it at least leads to an approximately unbiased estimator. However, it also becomes intuitively clear that variance reduction can be achieved by exploiting (10), for example by decoupling the estimation of $\tau(\theta)$ from $\frac{\partial\tau}{\partial\theta}(\theta)$, and generally by not settling for $\widehat{\frac{\partial\tau}{\partial\theta}}(\theta):=\frac{\partial\hat{\tau}}{\partial\theta}(\theta)$ as the estimator for $\frac{\partial\tau}{\partial\theta}(\theta)$.
3.1 Quantile Gradient Estimation
We can use Proposition 3.1 to design an algorithm that boosts the estimated quantile gradient. The idea is as follows: if we denote
$\eta(\theta)$ $:=\mathbb{E}\left[{\frac{\partial E}{\partial\theta}(\theta,X,Y)\,\big{|}\,A(\theta)}\right],$ $\Sigma(\theta)$ $:=\mathrm{cov}\left({\frac{\partial E}{\partial\theta}(\theta,X,Y)\,\big{|}\,A(\theta)}\right),$ (11)
$\eta_{\varepsilon}(\theta)$ $:=\mathbb{E}\left[{\frac{\partial E}{\partial\theta}(\theta,X,Y)\,\big{|}\,A_{\varepsilon}(\theta)}\right],$ $\Sigma_{\varepsilon}(\theta)$ $:=\mathrm{cov}\left({\frac{\partial E}{\partial\theta}(\theta,X,Y)\,\big{|}\,A_{\varepsilon}(\theta)}\right),$ (12)
for $\varepsilon>0$, where $A(\theta):=\{E_{\theta}(X,Y)=\tau(\theta)\}$ and $A_{\varepsilon}(\theta):=\{|E_{\theta}(X,Y)-\tau(\theta)|\leq\varepsilon\}$, then the terms in (11) and (12) are approximately equal if $\varepsilon\approx 0$. Subsequently, we can more efficiently estimate $\eta(\theta)=\frac{\partial\tau}{\partial\theta}(\theta)$, compared to the naive estimator $\hat{\eta}(\theta)=\frac{\partial\hat{\tau}}{\partial\theta}(\theta)$, by naive sample estimate of $\eta_{\varepsilon}(\theta)$. To this end, we propose the following $\varepsilon$-estimator
$\hat{\eta}(\theta):=\frac{1}{\sum_{i=1}^{n}1_{\hat{A}_{\varepsilon,i}(\theta)}}\sum_{i=1}^{n}1_{\hat{A}_{\varepsilon,i}(\theta)}\frac{\partial E}{\partial\theta}(\theta,X_{i},Y_{i}),$ (13)
from i.i.d. copies $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ of $(X,Y)$, where $\hat{A}_{\varepsilon,i}(\theta)=\{|E_{\theta}(X_{i},Y_{i})-\hat{\tau}(\theta)|\leq\varepsilon\}$.
Alternative gradient-boosting estimators can be constructed. Some examples include:
  • Ranking: sort the examples based on the distances $\{\big{|}E_{\theta}(X_{i},Y_{i})-\hat{\tau}(\theta)\big{|}\}_{i=1,\ldots,n}$, choose the “top” $m$ samples (smallest distances) for some suitably small $m$, and then set $\hat{\eta}(\theta)$ as the average of $\frac{\partial E}{\partial\theta}(\theta,X_{i},Y_{i})$ over those samples.
  • Kernel regression: consider some kernel $K_{h}(t)=\frac{1}{h}K\left(\frac{t}{h}\right)$ with $h>0$ for which $\int_{-\infty}^{+\infty}tK(t)\mathrm{d}t=0$, for instance $K(\cdot)=$ PDF of $\mathcal{N}(0,1)$ or $K(t)=\sigma(t)(1-\sigma(t)$ where $\sigma(\cdot)$ denotes the sigmoid function. Then, we can use the (modified) Nadaraya–Watson estimator $\hat{\eta}(\theta)=\frac{\sum_{i=1}^{n}K_{h}(E_{\theta}(X_{i},Y_{i})-\hat{\tau}(\theta))\frac{\partial E}{\partial\theta}(\theta,X_{i},Y_{i})}{\sum_{i=1}^{n}K_{h}(E_{\theta}(X_{i},Y_{i})-\hat{\tau}(\theta))}$.
  • Random splitting: split the $n$ examples into $m$ sub-sets of samples, apply some other estimator algorithm for $\eta(\theta)$ on each sub-set, and then set $\hat{\eta}(\theta)$ as the average of the individual estimates of $\eta(\theta)$.
Other closely related notions that could be adapted for the estimator $\hat{\eta}(\theta)$ and that could lead to reduced variance include importance sampling and smooth bootstrapping, but more generally, averages can be replaced by carefully constructed weighted sums. Particle filters could be feasible as a way of more efficiently estimate $\eta(\theta)$ by carrying over the previous estimates as $\theta$ is updated. It is also clear that various of these potential estimators are closely linked. For instance, if the threshold $\varepsilon>0$ is allowed to be sample-dependent, then the thresholding approach (13) and the ranking approach can be seen as equivalent by setting $\varepsilon=\inf\left\{\varepsilon^{\prime}>0:\sum_{i=1}^{n}1_{\hat{A}_{\varepsilon^{\prime},i}(\theta)}\geq m\right\}$.
3.2 Proposed Algorithm: VR-confTr
Suppose that a variance-reduced estimator for $\frac{\partial\tau}{\partial\theta}(\theta)$ has been already designed. Then, the new estimate for $\tau(\theta)$ and $\frac{\partial\tau}{\partial\theta}(\theta)$ can be plugged into expression (9) for the gradient of the conformal training risk function, before the expectations can be approximated by sample means, leading to the plug-in estimator for $\frac{\partial L}{\partial\theta}(\theta)$. Naturally, the plug-in gradient estimator is then passed through an optimizer in order to approximately solve (CRM). Our proposed pipeline, which we call variance-reduced conformal training (VR-ConfTr) algorithm, constitutes our main contribution and proposed solution to improve the sample inefficiency of ConfTr. The critical step of constructing the plug-in estimator is summarized in Algorithm 1. Additionally, the entire pipeline is illustrated in Figure 2.
Algorithm 1 Variance-reduced conformal training (VR-ConfTr)
0:   batch $B=\{(X_{1},Y_{1}),\ldots,(X_{2n},Y_{2n})\}$ of i.i.d. samples from $(X,Y)$, score function $E(\theta,x,y):\Theta\times\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$,conformal loss $\ell(\theta,x,y,\tau):\Theta\times\mathcal{X}\times\mathcal{Y}\times\mathbb{R}\to\mathbb{R}$,monotone transformation $\mathcal{F}:\mathbb{R}\to\mathbb{R}$,estimator $\hat{\tau}(\cdot)$ for $\tau(\theta)=Q_{\alpha}(E_{\theta}(X,Y))$,estimator $\widehat{\frac{\partial{\tau}}{\partial\theta}}(\cdot)$ for $\frac{\partial\tau}{\partial\theta}(\theta)$.
0:  output an estimate $\widehat{\frac{\partial L}{\partial\theta}}$ of the gradient $\frac{\partial L}{\partial\theta}(\theta)$ of the conformal training risk (ConfTr-risk)
1:  partition $B$ into $\{B_{\mathrm{cal}},B_{\mathrm{pred}}\}$, with $|B_{\mathrm{cal}}|=|B_{\mathrm{pred}}|=n$.
2:  $\hat{\tau}\leftarrow\hat{\tau}(B_{\mathrm{cal}})$ // estimate $\tau(\theta)$ using $B_{\mathrm{cal}}$
3:  $\widehat{\frac{\partial\tau}{\partial\theta}}\leftarrow\widehat{\frac{\partial\tau}{\partial\theta}}(B_{\mathrm{cal}})$ // estimate $\frac{\partial\tau}{\partial\theta}(\theta)$ using $B_{\mathrm{cal}}$
4:  ^$\hat{\ell}\leftarrow\frac{1}{|B_{\mathrm{pred}}|}\sum_{(x,y)\in B_{\mathrm{pred}}}\ell(\theta,x,y,\hat{\tau})$
5:  $\widehat{\frac{\partial\ell}{\partial\theta}}\leftarrow\frac{1}{|B_{\mathrm{pred}}|}\sum_{(x,y)\in B_{\mathrm{pred}}}\frac{\partial\ell}{\partial\theta}(\theta,x,y,\hat{\tau})$
6:  $\widehat{\frac{\partial\ell}{\partial\tau}}\leftarrow\frac{1}{|B_{\mathrm{pred}}|}\sum_{(x,y)\in B_{\mathrm{pred}}}\frac{\partial\ell}{\partial\tau}(\theta,x,y,\hat{\tau})$
7:  $\widehat{\frac{\partial L}{\partial\theta}}\leftarrow{{h}^{\prime}(\hat{\ell})}\left(\widehat{\frac{\partial\ell}{\partial\theta}}+\widehat{\frac{\partial\ell}{\partial\tau}}\widehat{\frac{\partial\tau}{\partial\theta}}\right)+\frac{\partial R}{\partial\theta}(\theta)$ // “plug-in” gradient estimator
8:  return  $\widehat{\frac{\partial L}{\partial\theta}}$
3.3 Theoretical Results
For the sake of concreteness, we focus our theoretical analysis on the thresholding approach (13). However, for simplicity and to avoid having to commit to any particular quantile estimator, we assume $\hat{\tau}(\theta)=\tau(\theta)$ in the analysis. Furthermore, we will assume $\varepsilon>0$ to be deterministic. Lastly, we will assume that in the event in which $\bigcup_{i=1}^{n}\hat{A}_{\varepsilon,i}(\theta)$ is empty, the estimator evaluates to $\hat{\eta}(\theta)=0$. With this, we can establish our main theoretical result in the following theorem.
Theorem 3.1 (Variance reduction).
Let $\hat{\eta}(\theta)$ be the estimator defined in (13) with $\hat{\tau}(\theta)=\tau(\theta)$. Then, the the bias and variance of the estimator can be characterized as follows:
$(i)$ $\hskip 19.91684pt\mathbb{E}\left[{\hat{\eta}(\theta)}\right]$ $=(1-q_{\varepsilon}(\theta)^{n})\eta_{\varepsilon}(\theta)$ (bias)
$(ii)$ $\hskip 14.22636pt\mathrm{cov}\left({\hat{\eta}(\theta)}\right)$ $\preceq\frac{2\Sigma_{\varepsilon}(\theta)}{p_{\varepsilon}(\theta)n}+\left[q_{\varepsilon}(\theta)\right]^{n}\eta_{\varepsilon}(\theta)\eta_{\varepsilon}^{\mathsf{T}}(\theta),$ (variance)
where $p_{\varepsilon}(\theta)=\mathbb{P}(A_{\varepsilon,i}(\theta))$ and $q_{\varepsilon}(\theta)=1-p_{\varepsilon}(\theta)$.
The main takeaway of result $(i)$ is that $\hat{\eta}(\theta)$ is an asymptotically unbiased estimator of $\eta_{\varepsilon}(\theta)$, but not $\eta(\theta)$. However, by definition we also have $\eta_{\varepsilon}(\theta)\approx\eta(\theta)$ for $\varepsilon\approx 0$. The second result $(ii)$, instead, shows that variance reduction is obtained by the proposed estimator, when compared to the naive estimator $\frac{\partial\hat{\tau}}{\partial\theta}(\theta)$. Further, for large $n$, the variance reduction is proportional to $p_{\varepsilon}(\theta)n$, which is equal to the (expected) proportion of samples that are ultimately used in the estimator. More precisely, the variance of the estimator is $\mathcal{O}\left(\frac{1}{p_{\varepsilon}(\theta)n}\right)$ as $\varepsilon\to 0$ or $n\to\infty$.
A key takeaway of $(i)$ and $(ii)$ is the explicit characterization of the bias-variance trade-off as a function of the threshold $\varepsilon>0$ and of the batch size $n$: for a given batch size $n$, a larger $\varepsilon$ increases the expected amount of samples used by the estimator, thus reducing its variance. However, larger $\varepsilon$ also increases the bias of the estimator towards the unconditional expectation $\mathbb{E}\left[{\frac{\partial E}{\partial\theta}(\theta,X,Y)}\right]$, where we make note that $\eta_{\varepsilon}(\theta)\to\eta(\theta)$ as $\varepsilon\to 0$.
4 Experiments
As a warm-up, we illustrate Theorem 3.1 on a synthetic Gaussian mixture model (GMM) dataset, depicted in Figure 3. We employ the $m$-ranking method with top $m=\frac{\alpha\,n}{\log\log n}$ samples. This ratio performs well across a variety of settings. As shown, our estimator (VR-ConfTr) reduces variance effectively, while the naive one (ConfTr) is sample inefficient.
Figure 3: Sample batch from GMM distribution (left) and corresponding bias and variance for the quantile gradient estimates (right).
4.1 Benchmark Datasets and Algorithms
We evaluate the effectiveness of VR-ConfTr against (i) a baseline model trained with standard cross-entropy loss (we refer to this method simply as Baseline), and (ii) the ConfTr algorithm proposed by Stutz et al. (2022). We perform experiments across benchmark datasets - MNIST Deng (2012), Fashion-MNIST Xiao et al. (2017a), Kuzushiji-MNIST Clanuwat et al. (2018) -, and a healthcare dataset comprising abdominal computed tomography scans, OrganAMNIST Yang et al. (2021). One of the main performance metrics that we consider is the length-efficiency of the conformal prediction sets produced by applying a standard CP procedure to the trained model. Other relevant metrics are the convergence speed and the variance across multiple runs. For the choice of the quantile gradient estimator $\frac{\widehat{\partial\tau}}{\partial\theta}$ of VR-ConfTr, we use the $m$-ranking approach presented in section 3.1. We choose this estimator because it is the more closely related to the one analyzed in Theorem 3.1. We investigate multiple possibilities for the choice of $m$, and more details on this tuning can be found in Appendix B. We provide extensive details about the training settings, the adopted model architectures, hyper-parameters and additional results in Appendix C.
In the next subsection, we present the summary of results obtained from evaluating the model after training. Given a number of epochs, we train a model over multiple runs. For the obtained model, we show: (i) the average accuracy and its standard deviation, and (ii), the average length efficiency and its standard deviation. In section 4.3, to illustrate further the improved training performance of VR-ConfTr over the original ConfTr algorithm and the variance reduction effect, we show the trajectories of relevant evaluation metrics - the conformal training loss defined in section 2, and the length efficiency - for all datasets and methods during training.
4.2 Summary of Evaluation Results
Table 1 presents the inefficiency results of the CP procedure applied post-training, and the accuracy of the trained model for each dataset, with the corresponding standard deviations.
Dataset Algorithm Avg Acc Std Acc Avg Size Std Size
MNIST Baseline 0.887 0.004 4.122 (+12%) 0.127
ConfTr Stutz et al. (2022) 0.842 0.141 3.990 (+8%) 0.730
VR-ConfTr (ours) 0.886 0.071 3.688 0.350
Fashion-MNIST Baseline 0.845 0.002 3.218 (+15%) 0.048
ConfTr Stutz et al. (2022) 0.799 0.065 3.048 (+9%) 0.201
VR-ConfTr (ours) 0.839 0.043 2.795 0.154
Kuzushiji-MNIST Baseline 0.872 0.046 4.982 (+6%) 0.530
ConfTr Stutz et al. (2022) 0.783 0.125 4.762 (+2%) 0.226
VR-ConfTr (ours) 0.835 0.098 4.657 0.680
OrganA-MNIST Baseline 0.552 0.017 4.823 (+2%) 0.748
ConfTr Stutz et al. (2022) 0.526 0.047 6.362 (+33%) 0.857
VR-ConfTr (ours) 0.547 0.021 4.776 1.178
Table 1: Summary of evaluation results. For VR-ConfTr, we show in percentage the average set size (Avg Size) improvement against ConfTr by Stutz et al. (2022). Third and fourth column are the average accuracy (Avg Acc) and its standard deviation (Std Acc).
The metrics reported in Table 1 are computed as averages over 5-10 training trials depending on the dataset. The way in which the number of the random training trials varies across the datasets is discussed in more detail in appendix C. Similarly to the approach followed by Stutz et al. (2022), we are mostly interested in the effectiveness of the different algorithms on the CP efficiency, and therefore we do not focus on improving the accuracy by using more advanced model architectures. To ensure a fair comparison for each dataset, we used the same exact model architecture across the three different methods (ConfTr, VR-ConfTr and Baseline). Furthermore, the training and evaluation hyper-parameters are identical across ConfTr and VR-ConfTr. For the CP procedure applied post-training, we use the standard THR method with $\alpha=0.01$. The average set-size for each method is reported over 10 different splits of the calibration and test data used for the conformal prediction procedure. The main takeaway from Table 1 is that VR-ConfTr improves over all considered metrics compared to ConfTr.
In terms of “length-efficiency”, VR-ConfTr is able to consistently achieve smaller prediction set sizes compared to both ConfTr and Baseline. It is important to note that the focus of our work is not to tune ConfTr to achieve better performance than Baseline, but rather to show that regardless of the performance of ConfTr and the hyper-parameters chosen, VR-ConfTr effectively provide performance improvements and training stability with the same hyper-parameters. Note that, similar to the results reported by Stutz et al. (2022), the Baseline architecture is sometimes able to achieve slightly higher accuracy than ConfTr and VR-ConfTr. It can be seen that VR-ConfTr consistently achieves higher accuracy compared to ConfTr. However, we stress that the objective of conformal training is to reduce the size of the prediction sets while preserving a similar accuracy as non-conformal training, and not to improve the accuracy.
4.3 On the Training Performance of VR-ConfTr
Here, we focus on the training performance of VR-ConfTr, with special attention to the speed in minimizing the conformal training loss described in section 2, and in minimizing the CP set sizes on test data. The results, which we illustrate plotting the evolution of the different metrics across epochs, validate the beneficial effect of the variance reduction technique and the superior performance of VR-ConfTr when compared to the competing ConfTr by Stutz et al. (2022).
Figure 4: Training curves for MNIST, Fashion-MNIST, Kuzushiji-MNIST, and OrganAMNIST. For each dataset, we show the training loss on top and corresponding test CP set sizes at the bottom at the end of each epoch, evaluated using the THR conformal predictor.
In Figure 4, we show the training performance for four datasets (MNIST, FMNIST, KMNIST and OrganAMNIST) illustrating two key metrics: (i) the evolution of the conformal training loss defined in section 2 and (ii) the test CP size across epochs. In the four plots on top, we show the comparison between the train loss evolution obtained using our VR-ConfTr against the one obtained by ConfTr. In the four plots at the bottom, we show the comparison between the test CP set sizes for VR-ConfTr, ConfTr and Baseline. In all the plots, we see that VR-ConfTr reaches smaller values of the loss and in significantly fewer epochs as compared to ConfTr. In the case of MNIST, for example, VR-ConfTr reaches a lower value of the loss in 10 times fewer epochs as compared to ConfTr. Similarly, for FMNIST VR-ConfTr achieves a smaller size in one third of epochs compared to ConfTr. For both Kuzushiji-MNIST and OrganA-MNIST, we notice that not only VR-ConfTr is faster, but it also gets to significantly smaller values of the loss. For the more challenging OrganA-MNIST dataset, this difference appears even more accentuated, not only in the training loss but also in the test CP set sizes. Notice that for all the three methods (VR-ConfTr, ConfTr and Baseline) we performed hyper-parameters tuning. Notably, in the case of the OrganA-MNIST dataset, we were not able to obtain an improvement with ConfTr in the final set size with respect to Baseline, which stresses the need for a method with improved gradient estimation, as the one we propose in this paper. More details on the grid-search over hyper-parameters and additional experiments for all algorithms can be found in appendix B and C.
5 Concluding Remarks and Future Directions
We formalized the concept of optimizing CP efficiency during training as the problem of conformal risk minimization (CRM). We identified a key source of sample inefficiency in the ConfTr method proposed by Stutz et al. (2022), which is a CRM method for length efficiency optimization. Our theoretical analysis elucidated the source of sample inefficiency, which lies in the estimation of the gradient of the population quantile. To address this issue, we introduced a novel technique that improves the gradient estimation of the population quantile of the conformity scores by provably reducing its variance. We show that, by incorporating this estimation technique in our proposed VR-ConfTr algorithm, the training becomes more stable and the post-training conformal predictor is often more efficient as well. Our work also opens up possibilities for future research in the area of CRM. Indeed, further methods for quantile gradient estimation could be developed and readily integrated with our “plug-in” algorithm, for which we can expect improved training performance.
Appendix A Proofs
In this appendix, we provide the proofs of all the theoretical results presented in the paper.
A.1 Proof of Lemma 3.1
Let $H(s)=\begin{cases}1,&s\geq 0\\ 0,&s<0\end{cases}$ denote the Heaviside step function, and let $H_{n}(s)=\Phi\left(\frac{s}{\sigma_{n}}\right)$ denote a smooth approximation, where $\Phi$ denotes the cumulative distribution function (CDF) of the standard Gaussian distribution, and $\sigma_{n}>0$ is a sequence such that $\sigma_{n}\to 0$ as $n\to\infty$. Note that $H_{n}(s)\to H(s)$ pointwise as $n\to\infty$, and that each $H_{n}$ is smooth.
By definition, we have
$\mathbb{P}\left[E(\theta,X,Y)\leq\tau(\theta)\right]=\alpha.$ (14)
which we can rewrite as
$\mathbb{E}\left[H\big{(}\tau(\theta)-E(\theta,X,Y)\big{)}\right]=\alpha.$ (15)
Since $0\leq H(s)\leq 1$ for all $s$, and $H_{n}(s)\to H(s)$ pointwise, by the Dominated Convergence Theorem, we have
$\alpha$ $=\mathbb{E}\left[H\big{(}\tau(\theta)-E(\theta,X,Y)\big{)}\right]$
$=\mathbb{E}\left[\lim_{n\to\infty}H_{n}\big{(}\tau(\theta)-E(\theta,X,Y)\big{)}\right]$
$=\lim_{n\to\infty}\mathbb{E}\left[H_{n}\big{(}\tau(\theta)-E(\theta,X,Y)\big{)}\right].$ (16)
Differentiating both sides with respect to $\theta$, we obtain
00 0 $=\frac{\partial}{\partial\theta}\lim_{n\to\infty}\mathbb{E}\left[H_{n}\left(\tau(\theta)-E(\theta,X,Y)\right)\right]$ (17)
$=\lim_{n\to\infty}\frac{\partial}{\partial\theta}\mathbb{E}\left[H_{n}\left(\tau(\theta)-E(\theta,X,Y)\right)\right]$ (18)
$=\lim_{n\to\infty}\mathbb{E}\left[\frac{\partial}{\partial\theta}H_{n}\left(\tau(\theta)-E(\theta,X,Y)\right)\right]$ (19)
$=\lim_{n\to\infty}\mathbb{E}\left[H_{n}^{\prime}\left(\tau(\theta)-E(\theta,X,Y)\right)\left(\frac{\partial\tau}{\partial\theta}(\theta)-\frac{\partial E}{\partial\theta}(\theta,X,Y)\right)\right],$ (20)
where $H_{n}^{\prime}$ denotes the derivative of $H_{n}$ with respect to its argument.
Justifications:
  • Equation 17: Differentiating both sides with respect to $\theta$ is valid because $\alpha$ is a constant independent of $\theta$.
  • Equation 18: to justify interchanging the limit and differentiation particularly
    $\frac{\partial}{\partial\theta}\lim_{n\to\infty}f_{n}(\theta)=\lim_{n\to\infty}\frac{\partial}{\partial\theta}f_{n}(\theta)$
    we need uniform convergence of $f_{n}(\theta)$ to $f(\theta)$ and uniform convergence of $\frac{\partial}{\partial\theta}f_{n}(\theta)$ to $\frac{\partial}{\partial\theta}f(\theta)$. Here,
    $f_{n}(\theta)=\mathbb{E}\left[H_{n}\left(\tau(\theta)-E(\theta,X,Y)\right)\right]$
    convergence pointwise to $\alpha$, which is constant, so their derivative converges to zero. the derivate of $\alpha$ with respect to $\theta$ is zero. so we can
    $0=\frac{\partial}{\partial\theta}\lim_{n\to\infty}\mathbb{E}\left[H_{n}\left(\tau(\theta)-E(\theta,X,Y)\right)\right]=\lim_{n\to\infty}\frac{\partial}{\partial\theta}\mathbb{E}\left[H_{n}\left(\tau(\theta)-E(\theta,X,Y)\right)\right]$
  • Equation (19): The interchange of differentiation and expectation is justified by the Leibniz Integral Rule, which allows to differentiate under the integral sign when the function inside the integral ($H_{n}$) and its derivative with respec to theta are continuous, and the integral of the absolute value of the derivative is finite. Here, $H_{n}(s)$ is inifinitely differentiable. $\tau(\theta)$ and $E(\theta,X,Y)$ are continuously differentibale with respect to $\theta$, and the derivative $\frac{\partial}{\partial\theta}H_{n}(\tau(\theta)-E(\theta,X,Y))$ is continuous in $\theta$ and integrable.
  • Equation (20): We apply the chain rule to differentiate $H_{n}\left(\tau(\theta)-E(\theta,X,Y)\right)$ with respect to $\theta$.
Define
$\delta_{n}(s)$ $=H_{n}^{\prime}(s)=\frac{1}{\sqrt{2\pi}\sigma_{n}}e^{-\frac{s^{2}}{2\sigma_{n}^{2}}},$
$\gamma_{n}(\theta,x,y)$ $=\delta_{n}\left(\tau(\theta)-E(\theta,x,y)\right),$
$\Delta(\theta,x,y)$ $=\frac{\partial\tau}{\partial\theta}(\theta)-\frac{\partial E}{\partial\theta}(\theta,x,y).$
Thus, we have
$\frac{\partial}{\partial\theta}\mathbb{E}\left[H_{n}\left(\tau(\theta)-E(\theta,X,Y)\right)\right]=\mathbb{E}\left[\gamma_{n}(\theta,X,Y)\Delta(\theta,X,Y)\right].$ (21)
let us denote $a_{n}\asymp b_{n}$ denote asymptotic equivalence, meaning that $\lim_{n\to\infty}\frac{a_{n}}{b_{n}}=1$, assuming $b_{n}\neq 0$. From equation (20), it follows that
$\mathbb{E}\left[\gamma_{n}(\theta,X,Y)\Delta(\theta,X,Y)\right]\asymp 0.$ (22)
Analyzing the Expectation:
Let $\varepsilon_{n}>0$ be any sequence such that $\varepsilon_{n}=o(\sigma_{n})$, meaning $\varepsilon_{n}/\sigma_{n}\to 0$ as $n\to\infty$. Define the set
$A_{\varepsilon_{n}}(\theta)=\left\{\omega\in\Omega:-\varepsilon_{n}<E(\theta,X(\omega),Y(\omega))-\tau(\theta)<\varepsilon_{n}\right\}.$ (23)
We can decompose the expectation in (22) as
$\mathbb{E}\left[\gamma_{n}\Delta\right]$ $=\mathbb{P}\left(A_{\varepsilon_{n}}(\theta)\right)\mathbb{E}\left[\gamma_{n}\Delta\,|\,A_{\varepsilon_{n}}(\theta)\right]+\mathbb{P}\left(A_{\varepsilon_{n}}^{c}(\theta)\right)\mathbb{E}\left[\gamma_{n}\Delta\,|\,A_{\varepsilon_{n}}^{c}(\theta)\right].$ (24)
Negligibility of the Second Term: On the complement $A_{\varepsilon_{n}}^{c}(\theta)$, the value $\tau(\theta)-E(\theta,X,Y)$ is either greater than $\varepsilon_{n}$ or less than $-\varepsilon_{n}$. Therefore, for $\ s\geq\varepsilon_{n}$ or $s\leq-\varepsilon_{n}$, $\delta_{n}(s)=H_{n}^{\prime}(s)$ becomes very small. Particularly
$\delta_{n}(s)=\dfrac{1}{\sqrt{2\pi}\,\sigma_{n}}\exp\left(-\dfrac{s^{2}}{2\sigma_{n}^{2}}\right)$
Since $\varepsilon_{n}=o(\sigma_{n})$ and $\sigma_{n}\to 0$, it follows that $\dfrac{\varepsilon_{n}}{\sigma_{n}}\to\infty$. For $|s|\geq\varepsilon_{n}$, we have: $\delta_{n}(s)\leq\dfrac{1}{\sqrt{2\pi}\,\sigma_{n}}\exp\left(-\dfrac{\varepsilon_{n}^{2}}{2\sigma_{n}^{2}}\right)$. from $\dfrac{\varepsilon_{n}}{\sigma_{n}}\to\infty$, it also follows that $\exp\left(-\dfrac{\varepsilon_{n}^{2}}{2\sigma_{n}^{2}}\right)\to 0$ and
$\gamma_{n}(\theta,X,Y)\leq\dfrac{1}{\sqrt{2\pi}\,\sigma_{n}}\exp\left(-\dfrac{\varepsilon_{n}^{2}}{2\sigma_{n}^{2}}\right)\to 0.$
Therefore, (24) becomes
$\mathbb{E}\left[\gamma_{n}\Delta\right]\asymp\mathbb{P}\left(A_{\varepsilon_{n}}(\theta)\right)\mathbb{E}\left[\gamma_{n}\Delta\,|\,A_{\varepsilon_{n}}(\theta)\right].$ (25)
Plugging in (25) into (22) we get:
$\mathbb{E}[\gamma_{n}(\theta,X,Y)\Delta(\theta,X,Y)\,|\,A_{\varepsilon_{n}}(\theta)]\asymp 0,$
By noting that $\mathbb{P}(A_{\varepsilon_{n}}(\theta))>0$ for all $n$ due to continuity of $x\mapsto E(\theta,x,y)$. We can then rewrite as
$\mathbb{E}[\gamma_{n}(\theta,X,Y),|\,A_{\varepsilon_{n}}(\theta)]\frac{\partial\tau}{\partial\theta}(\theta)\asymp\mathbb{E}\Bigg{[}\gamma_{n}(\theta,X,Y)\frac{\partial E}{\partial\theta}(\theta,X,Y)\,|\,A_{\varepsilon_{n}}(\theta)\Bigg{]}.$ (26)
When conditioned on $A_{\varepsilon_{n}}(\theta)$, we have $\gamma_{n}(\theta,X,Y)\asymp\delta_{n}(0)$. Indeed, note that $\delta_{n}(\varepsilon_{n})\leq\gamma_{n}(\theta,X,Y)\leq\delta_{n}(0)$, which we can rewrite as $\frac{\delta_{n}(\varepsilon_{n})}{\delta_{n}(0)}\leq\frac{\gamma_{n}(\theta,X,Y)}{\delta_{n}(0)}\leq 1$. Noting that $\delta_{n}(\varepsilon_{n})\asymp 1$, it indeed follows that $\gamma_{n}(\theta,X,Y)\asymp\delta_{n}(0)$. Therefore, (26) can be further simplified to
$\frac{\partial\tau}{\partial\theta}(\theta)\asymp\mathbb{E}\Bigg{[}\frac{\partial E}{\partial\theta}(\theta,X,Y)\,\big{|}\,A_{\varepsilon_{n}}(\theta)\Bigg{]}$ (27)
which readily leads to
$\frac{\partial\tau}{\partial\theta}(\theta)$ $=\lim_{n\to\infty}\mathbb{E}\Bigg{[}\frac{\partial E}{\partial\theta}(\theta,X,Y)\,\big{|}\,A_{\varepsilon_{n}}(\theta)\Bigg{]}$
$=\lim_{\varepsilon\to 0}\mathbb{E}\Bigg{[}\frac{\partial E}{\partial\theta}(\theta,X,Y)\,\big{|}\,A_{\varepsilon}(\theta)\Bigg{]}$
$=\lim_{\varepsilon\to 0}\mathbb{E}\Bigg{[}\frac{\partial E}{\partial\theta}(\theta,X,Y)\,\big{|}\,-\varepsilon<E_{\theta}(X,Y)-\tau(\theta)<\varepsilon\Bigg{]}$
$=\mathbb{E}\Bigg{[}\frac{\partial E}{\partial\theta}(\theta,X,Y)\,\big{|}\,E_{\theta}(X,Y)=\tau(\theta)\Bigg{]}.$
A.2 Proof of Theorem 3.1
We start with two preliminary results that we will use in the proof.
Some preliminaries. First, we recall a well-known result. Let $k\sim\text{Binomial}(n,p)$ be a random variable sampled from a Binomial distribution with $n$ trials and with probability $p$ of success. The following holds:
$\mathbb{E}\left[{\frac{1}{1+k}}\right]=\frac{(1-(1-p)^{n+1})}{(n+1)p}.$ (28)
Note that this follows from the following simple steps:
$\mathbb{E}\left[{\frac{1}{1+k}}\right]$ $=\sum_{k=0}^{n}\frac{1}{1+k}\cdot\binom{n}{k}p^{k}\left({1-p}\right)^{n-k}$
$=\frac{1}{p(n+1)}\sum_{k=0}^{n}\binom{n+1}{k+1}p^{k+1}(1-p)^{n-k}$
$=\frac{1}{p(n+1)}\sum_{j=1}^{n+1}\binom{n+1}{j}p^{j}(1-p)^{n+1-j}$
$=\frac{\left({1-(1-p)^{n+1}}\right)}{p(n+1)}.$
Next, we state another well-known identity. Let us consider the following recursion:
$a_{n+1}=\rho\,a_{n}+b,$
where $\rho>0$. Simply unrolling the recursion, we can obtain
$a_{n}=\rho^{n}a_{0}+b\left({\frac{1-\rho^{n}}{1-\rho}}\right)$ (29)
Proof of the Theorem.
Before we proceed, let us introduce some notation:
$G_{i}(\theta)$ $=\frac{\partial E}{\partial\theta}(\theta,X_{i},Y_{i})$ (30)
$A_{\varepsilon,i}(\theta)$ =$=\{\varepsilon\leq E(\theta,X_{i},Y_{i})-\tau(\theta)\leq\varepsilon\}$
$R_{\varepsilon,n}(\theta)$ $=\sum_{i=1}^{n}\chi_{A_{\varepsilon,i}(\theta)},$
$S_{\epsilon,n}(\theta)$ =$=\{i\in[n]:\chi_{A_{\epsilon,i}(\theta)}=1\}$
where $\chi_{A_{\varepsilon,i}(\theta)}$ is an indicator function for the event $A_{\varepsilon,i}(\theta)$, i.e.
$\chi_{A_{\varepsilon,i}(\theta)}=\begin{cases}1,&\textnormal{if }|E_{\theta}(X_{i},Y_{i})-\tau(\theta)|\leq\varepsilon\\ 0,&\textnormal{if }|E_{\theta}(X_{i},Y_{i})-\tau(\theta)|>\varepsilon.\end{cases}$ (31)
We are now ready to analyze the estimator $\hat{\eta}_{\varepsilon,n}(\theta)$ for ${\eta}_{\varepsilon}(\theta)$:
$\hat{\eta}_{\varepsilon,n}(\theta)=\begin{cases}\frac{1}{R_{\varepsilon,n}(\theta)}\sum_{i=1}^{n}\chi_{A_{\varepsilon,i}(\theta)}G_{i}(\theta),&\textrm{if}\hskip 2.84544ptR_{\varepsilon,n}(\theta)>0\\ 0&\textnormal{if }R_{\varepsilon,n}(\theta)=0\end{cases}$ (32)
where $\varepsilon$ and $n$ are denoted explicitly to remove ambiguity.
Equipped with the basic results established earlier in this subsection, we can proceed first with proving assertion (i). Note that, by definition (32), and because $\{X_{i},Y_{i}\}_{i=1}^{n}$ are sampled independently, we have
$\mathbb{E}\left[{\frac{1}{|S|}\sum_{i\in S}G_{i}|\bigcap_{i\in S}A_{\epsilon,i}(\theta)}\right]=\frac{1}{|S|}\sum_{i\in S}\mathbb{E}\left[{\chi_{A_{\epsilon,i}(\theta)}G_{i}(\theta)|A_{\epsilon,i}(\theta)}\right]=\eta_{\epsilon}.$ (33)
Also note that $S_{\epsilon,n}(\theta)=\emptyset$ is equivalent to $R_{\epsilon,n}(\theta)=0$, and that
$\mathbb{P}\left({S_{\epsilon,n}(\theta)=\emptyset}\right)=\mathbb{P}\left({R_{\epsilon,n}(\theta)=0}\right)=q^{n}$ (34)
with $q=1-p$ and $p=\mathbb{P}\left({A_{\epsilon,i}(\theta)}\right)$. Hence, we can get (i) as follows:
$\mathbb{E}\left[{\hat{\eta}_{\epsilon,n}(\theta)}\right]=q^{n}\mathbb{E}\left[{\hat{\eta}_{\epsilon,n}(\theta)|R_{\epsilon,n}(\theta)=0}\right]+(1-q^{n})\eta_{\epsilon},$ (35)
where we used the fact that $\sum_{S\neq\emptyset}\mathbb{P}\left({S_{\epsilon,n}(\theta)=S}\right)=1-\mathbb{P}\left({S_{\epsilon,n}(\theta)=\emptyset}\right)=1-q^{n}$.
Now, we prove (ii). We start by analyzing $\mathbb{E}\left[{\hat{\eta}_{\epsilon,n}(\theta)\hat{\eta}_{\epsilon,n}(\theta)^{\top}}\right]$:
$\mathbb{E}\left[{\hat{\eta}_{\epsilon,n}(\theta)\hat{\eta}_{\epsilon,n}(\theta)^{\top}}\right]$ $=\sum_{S\subseteq[n]}\mathbb{P}\left({S_{\epsilon,n}(\theta)=S}\right)\mathbb{E}\left[{\hat{\eta}_{\epsilon,n}(\theta)\hat{\eta}_{\epsilon,n}(\theta)^{\top}|S_{\epsilon,n}(\theta)=S}\right]$ (36)
$=\mathbb{P}\left({S_{\epsilon,n}(\theta)=\emptyset}\right)\mathbb{E}\left[{\hat{\eta}_{\epsilon,n}(\theta)\hat{\eta}_{\epsilon,n}(\theta)^{\top}|S_{\epsilon,n}(\theta)=S_{\epsilon,n}(\theta)}\right]$
$+\sum_{S\neq\emptyset}\mathbb{P}\left({S_{\epsilon,n}(\theta)=S}\right)\mathbb{E}\left[{\left({\frac{1}{|S|}\sum_{i\in S}G_{i}(\theta)}\right)\left({\frac{1}{|S|}\sum_{i\in S}G_{i}(\theta)}\right)^{\top}|\bigcap_{i\in S}A_{\epsilon,i}(\theta)}\right]$
$=\mathbb{P}\left({S_{\epsilon,n}(\theta)=\emptyset}\right)\mathbb{E}\left[{\hat{\eta}_{\epsilon,n}(\theta)\hat{\eta}_{\epsilon,n}(\theta)^{\top}|R_{\epsilon,n}(\theta)=0}\right]$
$+\sum_{S\neq\emptyset}\mathbb{P}\left({S_{\epsilon,n}(\theta)=S}\right)\frac{1}{|S|^{2}}\sum_{i\in S}\sum_{j\in S}\mathbb{E}\left[{G_{i}(\theta)G_{j}(\theta)^{\top}|A_{\epsilon,i}(\theta),A_{\epsilon,j}(\theta)}\right].$
Now note that
$\mathbb{E}\left[{G_{i}(\theta)G_{j}(\theta)^{\top}|A_{\epsilon,i}(\theta),A_{\epsilon,j}(\theta)}\right]$ $=\delta_{i,j}\mathbb{E}\left[{G_{i}(\theta)G_{i}(\theta)^{\top}|A_{\epsilon,i}(\theta)}\right]$ (37)
$+(1-\delta_{ij})\mathbb{E}\left[{G_{i}(\theta)|A_{\epsilon,i}(\theta)}\right]\mathbb{E}\left[{G_{j}(\theta)^{\top}|A_{\epsilon,j}(\theta)}\right]$
$=\mathbb{E}\left[{G_{i}(\theta)|A_{\epsilon,i}(\theta)}\right]\mathbb{E}\left[{G_{j}(\theta)^{\top}|A_{\epsilon,i}(\theta)}\right]$
$+\delta_{ij}(\mathbb{E}\left[{G_{i}(\theta)G_{i}(\theta)^{\top}|A_{\epsilon,i}(\theta)}\right]$
$-\mathbb{E}\left[{G_{i}(\theta)|A_{\epsilon,i}(\theta)}\right]\mathbb{E}\left[{G_{i}(\theta)^{\top}|A_{\epsilon,i}(\theta)}\right])$
$=\eta_{\epsilon}\eta_{\epsilon}^{\top}+\delta_{ij}\Sigma_{\epsilon}$
where we used the fact that $\{X_{i},Y_{i}\}_{i=1}^{n}$ are sampled i.i.d. and the definitions in (30). Now, we can proceed as follows:
$\mathbb{E}\left[{\hat{\eta}_{\epsilon,n}(\theta)\hat{\eta}_{\epsilon,n}(\theta)^{\top}}\right]$ $=q^{n}\mathbb{E}\left[{\hat{\eta}_{\epsilon,n}(\theta)\hat{\eta}_{\epsilon,n}(\theta)^{\top}|R_{\epsilon,n}(\theta)=0}\right]$ (38)
$+\sum_{S\neq\emptyset}\frac{\mathbb{P}\left({S_{\epsilon,n}(\theta)=S}\right)}{|S|}\left({|S|\eta_{\epsilon}\eta_{\epsilon}^{\top}+\Sigma_{\epsilon}}\right)$
$=q^{n}\mathbb{E}\left[{\hat{\eta}_{\epsilon,n}(\theta)\hat{\eta}_{\epsilon,n}(\theta)^{\top}|R_{\epsilon,n}(\theta)=0}\right]$
$+(1-q^{n})\eta_{\epsilon}\eta_{\epsilon}^{\top}+f_{n}\Sigma_{\epsilon}$
where we write
$f_{n}=\sum_{S\neq\emptyset}\frac{\mathbb{P}\left({S_{\epsilon,n}(\theta)=S}\right)}{|S|}.$ (39)
Now, we will show that
$f_{n}\leq\frac{2-p}{pn}.$ (40)
First, let us define the following function
$f(k)=\begin{cases}0,\ &\text{if }k=0\\ \frac{1}{k}&\text{if }k\geq 1\end{cases},$ (41)
and note that
$f_{n}=\mathbb{E}\left[{f(|S_{\epsilon,n}(\theta)|)}\right]=\mathbb{E}\left[{f(R_{\epsilon,n}(\theta))}\right].$ (42)
Now note that
$f_{n+1}$ $=\mathbb{E}\left[{f(R_{\epsilon,n+1}(\theta))}\right]$ (43)
$=\mathbb{P}\left({A_{\epsilon,i}(\theta)^{c}}\right)\mathbb{E}\left[{f(R_{\epsilon,n}(\theta))}\right]+\mathbb{P}\left({A_{\epsilon,i}(\theta)}\right)\mathbb{E}\left[{f(1+R_{\epsilon,n}(\theta))}\right]$
$=q\mathbb{E}\left[{f(R_{\epsilon,n}(\theta))}\right]+p\mathbb{E}\left[{f(1+R_{\epsilon,n}(\theta))}\right]$
$=qf_{n}+p\mathbb{E}\left[{\frac{1}{1+R_{\epsilon,n}(\theta)}}\right]$
$=qf_{n}+\frac{1-q^{n+1}}{n+1},$
where, in the last equation, we used the fact shown in the preliminaries (see (28)):
$\mathbb{E}\left[{\frac{1}{1+R_{\epsilon,n}(\theta)}}\right]=\frac{1-q^{n+1}}{p(n+1)}.$ (44)
Now let $a_{n}=nf_{n}$. We can write
$(n+1)f_{n+1}=(n+1)\left({qf_{n}+\frac{1-q^{n}}{n+1}}\right),$ (45)
from which we obtain the following recursion:
$a_{n+1}$ $=qnf_{n}+qf_{n}+(1-q^{n+1})$ (46)
$=qa_{n}+qf_{n}+(1-q^{n+1})$
$\leq qa_{n}+1+q,$
where we used the fact that $f_{n}\leq 1$ and that $1-q^{n}\leq 1$. With this recursion, we can now use the result illustrated in the preliminaries in (29) and get, using $q=1-p$,
$a_{n}\leq q^{n}a_{0}+\frac{1-q^{n}}{1-q}(1+q)\leq\frac{2-p}{p}.$ (47)
From the above inequality, we can conclude that
$0\leq f_{n}=\frac{a_{n}}{n}\leq\frac{2-p}{pn}.$ (48)
Pluggin this last result in (38), we can get
$\mathbb{E}\left[{\hat{\eta}_{\epsilon,n}(\theta)\hat{\eta}_{\epsilon,n}(\theta)^{\top}}\right]\leq q^{n}\mathbb{E}\left[{\hat{\eta}_{\epsilon,n}(\theta)\hat{\eta}_{\epsilon,n}(\theta)^{\top}|R_{\epsilon,n}(\theta)=0}\right]+(1-q^{n})\eta_{\epsilon}\eta_{\epsilon}^{\top}+\frac{2-p}{pn}\Sigma_{\epsilon}.$ (49)
We are now in the position to write and bound $\text{cov}(\hat{\eta}_{\epsilon,n}(\theta))$:
$\text{cov}\left({\hat{\eta}_{\epsilon,n}(\theta)}\right)$ $=\mathbb{E}\left[{\hat{\eta}_{\epsilon,n}(\theta)\hat{\eta}_{\epsilon,n}(\theta)^{\top}}\right]-\mathbb{E}\left[{\hat{\eta}_{\epsilon,n}(\theta)}\right]\mathbb{E}\left[{\hat{\eta}_{\epsilon,n}(\theta)^{\top}}\right]$ (50)
$\leq q^{n}\mathbb{E}\left[{\hat{\eta}_{\epsilon,n}(\theta)\hat{\eta}_{\epsilon,n}(\theta)^{\top}|R_{\epsilon,n}(\theta)=0}\right]+(1-q^{n})\eta_{\epsilon}\eta_{\epsilon}^{\top}+\frac{2-p}{pn}\Sigma_{\epsilon}$
$-\left({q^{n}\mathbb{E}\left[{\hat{\eta}_{\epsilon,n}(\theta)|R_{\epsilon,n}(\theta)=0}\right]+(1-q^{n})\eta_{\epsilon}}\right)$
$\cdot\left({q^{n}\mathbb{E}\left[{\hat{\eta}_{\epsilon,n}(\theta)|R_{\epsilon,n}(\theta)=0}\right]+(1-q^{n})\eta_{\epsilon}}\right)^{\top}$
$=\frac{2-p}{pn}\Sigma_{\epsilon}+q^{n}\eta_{\epsilon}\eta_{\epsilon}^{\top},$
where we used (i) and the fact that $\mathbb{E}\left[{\hat{\eta}_{\epsilon,n}(\theta)|R_{\epsilon,n}(\theta)=0}\right]=0$, which follows by (32).
Appendix B Additional experiments
In this section, we provide additional experimental results to complement the findings in the main paper
B.1 Additional Training Trajectories
Here, we present additional training trajectories, specifically the test loss and accuracy per epoch, for each dataset. These plots highlights the performance throughout the training process, providing further insights into convergence behavior and generalization performance. It can be seen that the test loss exhibits a pattern similar to the training loss in 4. In terms of accuracy, VR-ConfTr achieves higher accuracy than ConfTr.
Figure 5: Training curves for MNIST, Fashion-MNIST, Kuzushiji-MNIST, and OrganAMNIST. For each dataset, we show the test loss on the first row and tets accuracy on the bottom row at the end of each epoch.
B.2 Tuning VR-ConfTr: Number of points for Gradient Estimation (m)
In VR-ConfTr, the number of points ($m$) used to compute the gradient estimate plays a crucial role in the bias-variance trade-off. Consistent with the theory, increasing $m$ reduces the variance but potentially increases the bias of the gradient estimate. We conduct a grid search over the values $[4,6,8,10,16,20]$ for $m$ and report the results of tuning $m$ for MNIST, and Fashion MNIST, selecting the value of $m$ that experimentally provides the best trade-off between bias and variance. MNIST Results. As shown in Fig. 6, we observe a consistent reduction in the variance of gradient estimate as $m$ increases. However once we pass the optimal threshold the bias increases as can be seen by the higher values of the training loss as well as decrease in the size of the prediction sets. The figures corresponding to the loss on the training data per epoch, the loss on the test data per epoch, the accuracy evaluate on the test data per epoch, as well as the prediction set size evalauted on the test data per epoch.
Figure 6: Training trajectories for different values of $m$ on MNIST data
Fashion-MNIST. Similarly tuning $m$ on Fashion-MNIST shows that a value of $m=6$ provides the best results, as depicted in Fig. 7
Figure 7: Training trajectories for different values of $m$ on Fashion-MNIST data
B.3 Alternative Architecture
In this section, we compare the performance of VR-ConfTr on Kushuniji-MNIST using a simpler linear model architecture. The results indicate that regardless of the model architecture, the trends observed in terms of convergence speed and prediction set efficiency are consistent across datasets and architectures. Table 2 shows the average accuracy and set sizes for the two different models trained on K-MNIST.
Dataset Model Name Avg Accuracy Std Accuracy Avg Set Size Std Set Size
K-MNIST (Linear) Baseline 0.695 0.007 6.799 0.117
ConfTr 0.582 0.047 6.646 0.226
VR-ConfTr 0.612 0.033 6.488 0.148
K-MNIST (MLP) Baseline 0.872 0.046 4.982 0.530
ConfTr 0.783 0.125 4.762 0.226
VR-ConfTr 0.835 0.098 4.657 0.680
Table 2: Evaluation Results of KMNIST dataset trained with different model architectures
Appendix C Experimental Details
In this section we describe the experimental setup, including model architectures, dataset configurations, training protocol, testing procedure, and the corresponding hyper-parameters. The focus of the experiments is on evaluating the conformal prediction (CP) set sizes and ensuring a fair comparison between the baseline Conformal Training ConfTr and our proposed VR-ConfTr.
C.1 Dataset Configurations
We consider the benchmark datasets MNIST LeCun et al. (1998), Fashion-MNIST Xiao et al. (2017b), Kuzushiji-MNIST Clanuwat et al. (2018) and OrganAMNIST Yang et al. (2021).MNIST is a dataset of handwritten digits with 10 classes, and Fashion-MNIST consists of 10 fashion product categories. Kuzushiji-MNIST extends the MNIST paradigm by incorporating 10 classes of cursive Japanese characters. OrganAMNIST, derived from medical images, contains 11 classes of abdominal organ slices. The training, calibration, and testing splits for each dataset are summarized in Table 3. MNIST and Fashion-MNIST are provided by the torchvision library, while Kuzushiji-MNIST and OrganAMNIST are available from their respective repositories. For MNIST, Fashion-MNIST, and Kuzushiji-MNIST, 10% of the training set is reserved as calibration data. For OrganAMNIST, the validation set is used as the calibration data. During evaluation, we combine the calibration and test data and perform evaluations over 10 random splits of the combined dataset into calibration/test partitions. Model parameters are learned exclusively on the training data, while calibration and test data are used to evaluate the model as a black-box at the end of each epoch. The transformations applied to the dataset are as follows: for MNIST, Fashion-MNIST, and Kuzushiji-MNIST, images are normalized to have zero mean and unit variance, using a mean of 0.5 and a standard deviation of 0.5. For OrganAMNIST, images undergo random horizontal flips, random rotations of up to 15 degrees, and are normalized similarly.
Dataset Classes Image Size Training Set Calibration Set Test Set
MNIST 10 $28\times 28$ 55,000 5,000 10,000
Fashion-MNIST 10 $28\times 28$ 55,000 5,000 10,000
OrganMNIST 11 $28\times 28$ 34,561 6,491 17,778
Kuzushiji-MNIST 10 $28\times 28$ 55,000 5,000 10,000
Table 3: Dataset Splits
C.2 Model Architectures
In our experiments, we implemented all models using JAX Bradbury et al. (2018). We utilize a range of architectures including linear models, multi-layer perceptrons (MLPs), and modified ResNet architectures tailored for specific datasets. For the MNIST dataset, we employ a simple linear model, which consists of a single dense layer. The input images, reshaped from $28\times 28$ into a flattened vector of size $784$, are passed through a fully connected layer mapping the inputs directly to the $10$ output classes. This architecture provided a minimalistic baseline for comparison. For Fashion-MNIST, we use a multi-layer perceptron (MLP), with two hidden layers. We use 64 units per hidden layer, with ReLU activations Nair & Hinton (2010) , followed by a dense layer for the $10$ output classes. For Kuzushiji-MNIST, we utilize a similar MLP architecture. The model contains two hidden layers with 256 and 128 units, respectively. The input data is flattened and passed through these fully connected layers with ReLU activations. For OrganAMNIST, we used a residual network, inspired by the ResNet architecture from He et al. (2016) , with modifications. The model consists of an initial convolutional layer followed by four stages of residual blocks, each with two layers. Each residual block uses $3\times 3$ convolutions with ReLU activations. The number of output channels doubles after each state $(64,128,256,512).$ Global average pooling is applied before the final fully connected layer, which maps the pooled feature representations to the 11 output classes. We do not attempt to optimize the model architectures in order to solve the datasets with high accuracy. Instead, we focus on the conformal prediction results, and ensure that the architecture used across different algorithms are identical for a fair comparison.
C.3 Training Details
Similar to Stutz et al. (2022), we trained all models using Stochastic Gradient Descent (SGD) with Nesterov momentum Sutskever et al. (2013). The learning rate follows a multi-step schedule where the initial learning rate was decreased by a factor of 0.1 after 2/5, 3/5, and 4/5 of the total number of epochs. The models were trained using cross-entropy-loss for Baseline training, and for ConfTr and VR-ConfTr based on the size-loss as described by Stutz et al. (2022). During training, we set the conformal prediction threshold parameter $\alpha=0.01$. To ensure statistical robustness, we conducted multiple randomized training trials for each dataset, using a different random seed for each trial. Specifically, we performed 10 training trials for MNIST and 5 training trials each for FMNIST, KMNIST, and OrganAMNIST. During each trial, a unique random seed was used to initialize the model and optimizer, ensuring that each trial followed a distinct learning trajectory. The corresponding training trajectories, i.e the training loss, testing loss, accuracy and CP set sizes evaluated on the test data at the end of every epoch, were averaged over these randomized trials to provide a smooth and general view of the model’s performance. The key hyper-parameters used for training are listed in Table 4. These hyper-parameters include size weight which scales the loss term associated with the size of the CP sets during training, alpha $\alpha$ corresponding to the miscoverage rate is set to 0.01. batch size for SGD, learning rate for the optimizer, and the number of epochs for which the model is trained for.
Hyper-parameter MNIST Fashion-MNIST Kuzushiji-MNIST OrganA-MNIST
Batch Size 500 500 500 500
Training Epochs 50 150 100 100
Learning Rate 0.05 0.01 0.01 0.01
Optimizer SGD SGD SGD SGD
Temperature 0.5 0.1 0.1 0.5
Target Set Size 1 0 1 1
Regularizer Weight 0.0005 0.0005 0.0005 0.0005
Size Weight 0.01 0.01 0.01 0.1
Alpha ($\alpha$) 0.01 0.01 0.01 0.01
Num. of Pts for Gradient 6 6 4 4
Table 4: Training and evaluation Hyper-parameters for each dataset.
C.4 Evaluation Details
The evaluation of our models was conducted in two stages: (1) computing the test accuracy for each model after training, and (2) evaluating the conformal prediction (CP) set sizes and coverage over multiple test splits. The goal was to ensure both accuracy and conformal prediction performance are consistently reported across randomized trials and test splits. Test Accuracy: For each dataset, the test accuracy of the trained models was evaluated on the test data, and the results were averaged over the randomized training trials. CP set sizes To compute the average conformal prediction (CP) set size, we first combine the holdout calibration and test data. We then randomly split this combined data into calibration and test portions, repeating the process 10 times. For each split, we apply the CP THR algorithm with $\alpha=0.01$ to compute the prediction set sizes on the test portion, and the results are averaged across the 10 random splits. The cardinality of each split is consistent with the dataset configurations outlined in Table 3. This procedure is performed for each trained model, and the final reported results are averaged across both the training trials and testing splits.
C.5 Differences from ConfTr reports
We report the performance of Conftr with a batch size of 100 for Fashion-MNIST, as originally reported by Stutz et al. (2022), selected for optimal performance. While a batch size of 500 yields smaller set sizes, it results in a slight ( 1%) decrease in accuracy. For completeness, we include the results for both configurations.
Model Batch Size Average Accuracy Std Accuracy Average Set Size Std Set Size
ConfTr 100 0.809 0.051 3.125 0.197
ConfTr 500 0.799 0.065 3.048 0.201
VR-ConfTr 500 0.839 0.043 2.795 0.154
Table 5: Final Evaluation Results for Fashion-MNIST
Retrieving exact reported set sizes as Stutz et al. (2022): Our experimental results and trends align with those reported in Stutz et al. (2022). However, the smaller set sizes for Conftr on MNIST and FMNIST in their paper are likely due to their use of more advanced/different architectures. Despite this, the overall trends— Conftr outperforming Baseline, and VR-Conftr outperforming Conftr—remain consistent regardless of the model. Our focus is on a fair comparison across algorithms by using the same architecture, rather than reproducing the exact figures or architectures from Stutz et al. (2022).